The work function of sodium metal is 2.3 eV. What is the longest wavelength of light that cause

photoelectric emission from sodium


1
Expert's answer
2021-09-13T02:42:10-0400

Work function W = 2.3 eV

"1\; eV = 1.6 \times 10^{-19} \;J \\

W = 2.3 \times 1.6 \times 10^{-19} = 3.68 \times 10^{-19} \;J"

longest wavelength "\u03bb_0" will be

"W= \frac{hc}{λ_0} \\

λ_0 = \frac{hc}{W} \\

h=6.626 \times 10^{-34} \;Js \\

c= 3 \times 10^8 \;m/s \\

λ_0 = \frac{6.626 \times 10^{-34} \times 3 \times 10^8}{3.68 \times 10^{-19}} \\

= 5.4 \times 10^{-7} \;m"